The resistance (Rn)of an electrical circuit, containing resistors R1,R2,and R3 in parallel is given by


1
Expert's answer
2021-10-11T01:38:13-0400
"\dfrac{1}{R_n}=\dfrac{1}{R_1}+\dfrac{1}{R_2}+\dfrac{1}{R_3}"

"\dfrac{1}{R_n}=\dfrac{R_1R_2+R_1R_3+R_2R_3}{R_1R_2R_3}"

"R_n=\dfrac{R_1R_2R_3}{R_1R_2+R_1R_3+R_2R_3}"